solve the differential equation to the indicated inintial conditions; "y+4y=0" "y(0)=4" "y(0)=6"


1
Expert's answer
2021-10-20T18:28:03-0400

Here, question is incorrect as "y+4y=0" is not a differential equation also "y(0)=4, y(0)=6"

is not possible as we are getting two different "y" values at "x=0."

Let us take the differential equation as "y'+4y=0" and the boundary condition as "y(0)=4."

"y'+4y=0
\\\Rightarrow \frac{dy}{dx}+4y=0\\
\Rightarrow \frac{dy}{dx}=-4y\\\Rightarrow \frac{dy}{y}=-4dx"

Integrating both sides, we get

"\int \frac{dy}{y}=\int(-4)dx
\\\Rightarrow ln \ y = -4x+c \ ...(1)"

Now, using boundary condition as y(0)=4

"ln \ 4=-4\times0+c\Rightarrow c=ln \ 4"

From equation (1), we get

"ln \ y = -4x+ln \ 4
\\\Rightarrow 4x=ln\ 4 - ln\ y
\\\Rightarrow 4x=ln\ (\frac{4}{y})
\\\Rightarrow e^{4x}=\frac{4}{y}
\\\Rightarrow y.e^{4x}=4"